If you have 14% milk fat ice cream base and 10% milk fat ice cream base, what combination amounts of the two would make 1000 ml of 12% ice cream base?
Answer by stanbon(75887)

If you have 14% milk fat ice cream base and 10% milk fat ice cream base, what combination amounts of the two would make 1000 ml of 12% ice cream base?
-------------------
Equation:
fat + fat = fat
0.14x + 0.10(1000-x) = 0.12*1000
---
Multiply thru by 100 to get:
14x + 10*1000-10x = 12*1000
4x = 2*1000
x = 500 ml (amt. of 14% solution needed)
========================================
Cheers,
Stan H.
============